# Axion superradiance

### v1

@@ Abstract
Light bosonic fields may suffer an instability around a rotating compact object. This process, known as superradiance, leads to the exponential amplification of the field around a black hole or neutron star, while the spin of the central object is correspondingly depleted. The discovery of a highly spinning black hole could therefore be used to constrain the existence of light bosons such as axions in a particular range of masses. These constraints apply for very low non-gravitational couplings between the boson and the Standard Model, offering a powerful search strategy for new physics. However, care must be taken to include the more complex effects of the black hole’s astrophysical environment. Conversely, stellar superradiance could allow us to probe additional non-gravitational interactions between a new boson at the stellar matter. In this article, I will discuss the current status and future directions of axion superradiance. This is a contribution to the proceedings of the 3rd General Meeting of the COST Action COSMIC WISPers.
@@ 2 Black hole superradiance
  • A superradiant instability exists when $\omega_{R}<m\Omega$, where $m$ is the azimuthal qauntum number of the bound state and $\Omega$ is the angular velocity of the BH.
@@
These results tell us that efficient superradiance is not possible for Standard Model bosons. The electroweak bosons are much too heavy for observable superradiance with either astrophysical or supermassive BHs. Detailed study of BH superradiance involving a photon with a plasma mass suggests that this is also probably not observable [4]. However, superradiance may be a highly effective probe of light Beyond the Standard model bosons such as axions and dark photons.

### v2

@@ Abstract
Light bosonic fields may suffer an instability around a rotating compact object. This process, known as superradiance, leads to the exponential amplification of the field around a black hole or neutron star, while the spin of the central object is correspondingly depleted. The discovery of a highly spinning black hole could therefore be used to constrain the existence of light bosons such as axions in a particular range of masses. These constraints apply for very low non-gravitational couplings between the boson and the Standard Model, offering a powerful search strategy for new physics. However, care must be taken to include the more complex effects of the black hole’s astrophysical environment. Conversely, stellar superradiance could allow us to probe additional non-gravitational interactions between a new boson at the stellar matter. In this article, I will discuss the current status and future directions of axion superradiance. This is a contribution to the proceedings of the 3rd General Meeting of the COST Action COSMIC WSIPers.
@@ 2 Black hole superradiance
  • A superradiant instability exists when $\omega_{R}<m\Omega$, where $m$ is the azimuthal quantum number of the bound state and $\Omega$ is the angular velocity of the BH.
@@
These results tell us that efficient superradiance is not possible for Standard Model bosons. The electroweak bosons are much too heavy for observable superradiance with either astrophysical or supermassive BHs. Detailed study of BH superradiance involving a photon with a plasma mass suggests that this is also probably not observable [4]. However, superradiance may be a highly effective probe of light Beyond the Standard model bosons such as axions and dark photons, as first noted in [5].